OpenStudy (anonymous):

One of the lines in the emission spectrum of sodium has a wavelength of 590 nm what is the frequency of the line?

OpenStudy (anonymous):

my teacher Mr. burns changed 590 to 590x10 with the exponent of -9 m

OpenStudy (dan815):

yes because 590 NANOmeters is 590*10^-9 meters

OpenStudy (anonymous):

ok.

OpenStudy (unklerhaukus):

\[c=f\lambda\] \[f = c/\lambda\\ \quad= \frac{3.00\times10^8\,\text m/\text s}{590\times10^{-9} \,\text m}\\ \quad=\frac{3.00}{590}\times10^{8-(-9)}\,/\text s\\ \quad=\]